Circular No. 6265

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

α MONOCEROTID METEORS 1995 Numerous reports have been received, from several locations in Eu-

rope, of an excellent display of this occasional, short-lived meteor shower around Nov. 22.06 UT. Mentioned by Olivier (1936, Pop. Astron. 44, 88) as perhaps having a 10-year periodicity, and discussed at length by Kresák (1958, Bull. Astron. Inst. Czech. 9, 88), the shower was widely anticipated to recur in 1995, Jenniskens (1995, WGN 23, No. 3, 84) predicting the brief maximum as occurring some time during Nov. 22.00–22.25 UT. P. Spurný and J. Borovicka, Ondřejov Observatory, report monitoring during Nov. 21.96–22.11, Monocerotids first being detected at 22.050, with more than six during each 2-min interval until 22.067 and some continuing activity to 22.098; during one minute at 22.058 a maximum of seven meteors was observed. A. Gomez, Instituto de Astrofisica de Canarias, reports observations from Almeria by L. Bellot, A. Roman and F. Reyes of as many as 70 Monocerotids during 22.055–22.071; the limiting visual magnitude was ∼ 6.2, and most of the meteors were in the magnitude range 0–2. J. Rendtel, International Meteor Organization, Potsdam, noted 34 shower meteors from a radiant of α = 113^{o}, δ = −3^{o} during 22.042–22.076 (limiting magnitude 6.15), yielding an average ZHR (assuming a population index of r = 2.5) of 90 (with an equivalent ZHR of up to 190 during 10 min centered on 22.062; on combining his results with those of S. Molau in Chemnitz, he revised this to 220 ± 50). G. Forti, Arcetri Observatory, reports observations by R. Haver and R. Gorelli from Frasso Sabino, indicating a radiant of α = 112^{o}.5, δ = −3^{o} and a peak within 5 min of 22.059, there being a sharp rise and fall and no meteors observed 15 min before or after the peak; the meteors were yellow and blue, and few were brighter than mag 0. Z. A. Nagy, K. Sárenczky and I. Tepliczky, observing from Vértes Mountain (Hungary), collectively recorded Monocerotids during 22.045–22.087, counting 139 shower members during 22.049–22.072 (with maximum activity during 22.060–22.065), 90 percent of them from a compact radiant at α = 116^{o}, δ = +4^{o}, and the others from a diffuse radiant between α CMi and δ Mon; several meteors were around mag 0. J. Gerboš, P. Rapavý and V. Hrušovský, Rimavská Sobota (Slovakia), collectively registered more than 600 Monocerotids during one hour beginning 22.049, the maximum activity in 5 min being centered on 22.058. C. Steyaert reports that M. De Meyère, monitoring in Deurle (Belgium) a 100-kW radio station in Budapest, 1170 km away, at 66.51 MHz, detected a significant (fourfold) enhancement in forward-scattering rates during the hour beginning 22.042, the number of meteor reflections lasting at least 0.027 second being 232.

1995 November 22 Brian G. Marsden